OpenStudy (anonymous):

-6+3x=-9 WILL GIVE MEDAL FOR CORRECT ANSWER!

OpenStudy (dan815):

will u give medal if i show you how to do it too

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

-6+3x=-9 +6 +6 3x=-3x dive 3 from each side and get x=-1

OpenStudy (dan815):

lol

OpenStudy (anonymous):

divide*